OpenStudy (anonymous):

solve in interval notation: x^4-4x^2>0 Help please!!!

OpenStudy (anonymous):

\[x^{2}(x ^{2}-4). \] And then do a sign chart to find that the answer is: \[(-\infty,2) and (2,\infty)\]

OpenStudy (anonymous):

do you mean -\[(-\infty,-2) U (2,\infty)\] cause when I tested the middle I got that it was a no

OpenStudy (anonymous):

x^2- 4 is the plus and minus or is that x^2+4 ?